OpenStudy (anonymous):

Write an equation for the statement. Half the number y is equal to y minus ten. (Points : 1)

OpenStudy (anonymous):

hello people can you help :)

OpenStudy (anonymous):

Would it be 1/2y=y-10? Do you have multiple choice answers?

OpenStudy (anonymous):

Yeah i get the same. 1/2y=y-10